Larson Manufacturing is considering purchasing a new injection-molding machine for $250,000 to expand its
production capacity. It will cost an additional $20,000 to do the site installed. With the new injection-molding
machine installed, Larson Manufacturing expects to increase its revenue by $90,000 per year. The machine will
be used for five years, with an expected salvage value of $75,000.
a. Compute the NPV of the project at an interest rate of 12% .
b. Based on NPV, would the purchase new injection-molding machine be justified?
Answer #1

a)

Initial Cost=Co=Purchase Cost+Installation cost=250000+20000=$270,000

Increase in Revenue per year=R=$90,000

Interest Rate=i=12%

Useful life=n=5 years

Salvage Value=S=$75000

Increase in revenue is same as ordinary annuity of $90000 for 5 years. So, PV of increase in revenues is equal to PV of ordinary annuity of $90000 for 5 years @12% interest rate.

PV of salvage is equal to PV of $75000 to be received at the end of year 5.

NPV of project is given by

NPV of project=-Co+R*(P/A,12%,5)+S*(P/F,12%,5)

Let us calculate interest factors.

(P/F,0.12,5)=1/(1+0.12)5=0.567427

NPV of project=-270000+90000*3.604776+75000*0.567427=$96,986.87

b)

We find that NPV of project is positive value. It means that project adds to company's present worth. Purchase is justified in this case.